OpenStudy (anonymous):

name the property for.. 0+x=x 6+(-6)=0

OpenStudy (anonymous):

\[8\left[ -10/4 \right]=\left[ -10/4 \right]8\]

OpenStudy (turingtest):

learn this list http://www.capitan.k12.nm.us/teachers/shearerk/basic_rules_of_algebra.htm all the answers are on there, and if you learn this it will serve you for the rest of all math that you study